OpenStudy (anonymous):

use rational exponents to simplify ^4√x16

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\sqrt[4]{16x}\]?

OpenStudy (anonymous):

if so, \[\sqrt[4]{16x}=2\sqrt[4]{x}=2x ^{1/4}\]